OpenStudy (anonymous):

A sledgehammer weighs 18 pounds on Earth and 4 pounds on the planet Namar. Based on the numbers, an 80–pound earthling would weigh about _____ pounds on Namar.

OpenStudy (texaschic101):

18/4 = 80/x (18 lbs on Earth to 4 lbs on Namar = 80 lbs on earth to x lbs on Namar) cross multiply 18x = (4)(80) 18x = 320 x = 320/18 x = 17.77 rounds to 18 lbs on Namar

OpenStudy (anonymous):

oh thxz

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

were u frm

OpenStudy (texaschic101):

Texas :)

OpenStudy (anonymous):

were

OpenStudy (anonymous):

360